How much water must be added to 100ml of 50 % acid/water solution to obtain a 30% solution.
Answer by Nate(3500)

x amount of 0% acid (pure H2O)
(100(50%) + x(0%))/(100 + x) = 30%
100(0.5) + 0 = 0.3(100 + x)
50 = 30 + 0.3x
20 = 0.3x
200/3 = 66.6666ml = x